# Design Space Exploration of Embedded SoC Architectures for Real-Time Optimal Control

### v1

@@ Abstract
Empowering resource-limited robots to execute computationally intensive tasks such as package delivery, high-precision surgical assistance, and autonomous drone swarming is challenging. The diversity of algorithms and kernel types in these workloads causes bottlenecks in different systems, depending on application requirements, making a single hardware architecture inadequate for all robotic applications. This project provides a comprehensive design space exploration to determine optimal hardware computation architectures suitable for model-based control algorithms. We profile and optimize representative architectural designs across general-purpose scalar and vector processors and specialized accelerators. Specifically, we compare CPUs, vector machines, and domain-specialized accelerators with kernel-level benchmarks and end-to-end representative robotic workloads. Our exploration provides a quantitative performance, area, and utilization comparison and analyzes the trade-offs between these representative distinct architectural designs. We demonstrate that the variation of hardware architecture choices depends on workload characteristics and application requirements. Finally, we explore how architectural modifications, software and system optimization can alleviate bottlenecks and enhance utilization.
@@ I Introduction
Robotic systems such as humanoids, legged robots, unmanned aerial vehicles, and autonomous vehicles are widely deployed across diverse sectors, demonstrating their potential to revolutionize our daily lives. Robust model- and learning-based algorithms enable robots to efficiently carry out complex real-world tasks such as search and rescue, human-robot interaction, and package delivery. With the rapid evolution of robotic applications, the computational demands of end-to-end robotic pipelines are escalating dramatically. Besides, the
However, the computational complexity of those applications, coupled with real-time requirements, poses significant challenges to the compute platform on robotic devices. Many robotic platforms have limited computing power on board, especially micro/nano-scaled robots, which can only support microcontrollers or embedded SoCs.
Most microcontrollers and embedded SoCs either lack the computational power to meet algorithm latency requirements or consume excessive energy, rendering them impractical for battery-powered edge devices. As robots become more complex and miniaturized, onboard resources, such as compute capabilities, memory storage, and bandwidth, become increasingly constrained. The weight, power constraints, and area limitations restrict the ability of robotic systems to carry out broad real-life tasks in latency-critical or power-limited scenarios. Characterizing the performance of hardware architectures is critical to designing robotic systems, as the design and performance of a robotic SoC directly impacts the end-to-end performance and behavior of the robot [18].
Given the complexity and diversity of algorithms and kernel types required by modern robotic applications, relying on a single hardware architecture to meet these demands is becoming increasingly impractical. Observing the heterogeneity in modern SoCs, this project aims to explore the computational architectures that best address the increasing demands of model-based control algorithms. We conduct a comprehensive design space exploration to identify the most effective computational platforms and architectures capable of meeting these intensifying demands. In this paper, we categorize various compute platforms in Section II, evaluate the computational and resource requirements of robotic workloads in Section IV, and present detailed quantitative profiling of different architectural designs applied to these workloads in Section VI. This exploration is crucial for enhancing the efficiency, scalability, and real-time performance of robotic systems operating in increasingly resource-constrained environments.
@@ II Overview of Architectures
CPUs are the most commonly used general-purpose computing platforms. CPUs, as general processors, are designed to handle a broad range of kernels. Due to this attribute, cheap CPUs, especially microcontrollers, were traditionally used for simple rule-based robotic kernels such as PID control solvers. However, microcontrollers fail to meet the computational demand for real-time robotic tasks with scaled-up models and DNN-based algorithms. This is mainly due to microcontrollers being typically slow with a compute frequency of 200mHZ. One solution to maximize a CPU’s available compute is by exploiting the out-of-order and superscalar architecture to drive instruction-level parallelism. However, the performance margin comes with the cost of area overhead and power consumption. An out-of-order and superscalar CPU can achieve up to 100 GFLOP/Sec. However, its power efficiency is less than 1 GOP/J, which is usually not compatible with power-constrained embedded devices [26].
@@
Vector extensions allow lightweight front-end CPUs to achieve similar performance as out-of-order and superscalar CPUs, while dramatically simplifying implementation by eliminating excess logic and simplifying implementation. Vector machines exploit data parallelism by processing multiple data elements with a single instruction without adding too much power and area overhead. Various vector architectures range from traditional long-vector machines, as presented in Cray-style vector machines [21] to Packed SIMD ISA (P-SIMD) machines that utilize VLIW instruction encodings, [12], to contemporary novel short-vector machines [28]. These architectures adeptly support a broad array of operations, such as scalar multiplication, general vector multiplication (GEMV), and matrix-matrix multiplication (GEMM). While vector cores are highly flexible and can handle a wide range of computations, they may not be as efficient for GEMM as dedicated systolic arrays or other matrix-specific accelerators, especially in handling large, batched operations. This is because vector machines scale inadequately with a large datapath compared to a systolic array-style processor. As datapath (DLEN) scales up to feed in data sufficiently to the increased number of lanes, the bandwidth of the register file is also required to scale up; the scaling factor is equivalent to the bandwidth of DLEN.
@@
GPUs are designed to drive data-level and thread-level parallelism with up to over 100 thousand cores running simultaneously, enabling massive parallelism. Only a few robotic workloads[5, 15] can be benefited from such massive parallelism. However, a typical GPU can reach up to 10 TOPS/Sec of performance. Its high throughput makes it suitable for tasks that involve heavy matrix computations and graphics rendering, and other non-real-time robotic workloads. GPUs also exhibit Thread-Level Parallelism (TLP) by running thousands of threads simultaneously to handle extensive graphical calculations and data processing tasks efficiently. However, a conventional GPU’s die area is typically around 600 mm² to over 800 mm² [7], which consumes too much area to be placed on micro- or nano-robotics where the system has limited space on a chip and faces weight constraints. In terms of energy consumption, GPUs consume 100W 350W of power[27], which is far beyond the power budget of most resource-limited robotic systems.
@@
In addition to the general-purpose processors, ASICs and FPGAs have recently been proposed as feasible platforms for power-constraint robotic applications since they offer higher energy efficiency than CPUs and GPUs. Both ASICs and FPGAs consume relatively less power and are often integrated into small systems with limited memory. Featured by the high reconfigurability and programmability, FPGAs adapt well to the involving robotic algorithms with relatively lower margins than ASICs. Partial Reconfiguration (PR) extends this flexibility, allowing dynamic reconfiguration of a portion of the FPGA at runtime[24]. However, the highly specialized programming model, challenges in mapping irregular algorithms with control flow, and high resource overhead from configuration/routing logic make maximizing the utilization of all processing elements (PEs) challenging.
@@
With the integration of domain-specific accelerators, ASICs and FPGAs can parallelize computations while leveraging the specific characteristics of the kernels to eliminate excess logic and simplify implementation[24] and offload specific computational tasks to the accelerators. Research [25, 13, 16] has been devoted to developing hardware-aware algorithms for FPGAs and has been proposed to achieve software-equivalent performance, making ASICs and FPGAs capable of meeting real-time requirements [14, 22, 23].
@@
Particularly, spatial array style domain-specific accelerators have been attributed to their energy efficiency and compute throughput in compute-intensive workloads such as neural networks. A typical spatial array accelerator architecture can achieve an efficiency of approximately 34.4 GOP/J while consuming 278 mW at a frame rate of 35 FPS for the AlexNet model[4]. The benefit of re-purposing DNN accelerators for robot computing is based on the observation that neural network accelerators are optimized for matrix linear algebra [9, 11] while major robotic algorithms also mainly consist of DNN workloads and small-sized general matrix-matrix multiplication (GEMM) and vector multiplication(GEMV) [8, 2]. Most robotics chips already contain an ASIC neural net inference accelerator to speed up DNN-based workloads like perception, vision, and online learning. Suppose it can be shown that the same accelerator is also capable of speeding up robot dynamics computations. In that case, the overall SoC architecture can benefit from significant savings in the chip area, make room for other specialized units, or instead enable a physical design that is more cost-effective for ASIC implementation.
@@
Besides DNN accelerators, research efforts[22, 25, 13, 5, 15, 16, 14, 23, 19] have been dedicated to design domain-specific robotic accelerators exclusively for robotic algorithms, focusing on a particular kernel defined for optimization problems for specific types of robots. Although domain-specific accelerators can offer improved efficiency over general-purpose processors on specific robotics workloads, the dramatic amount of engineering cost due to its specific programmable interface with its dedicated domain-specific language (DSL); Besides, due to the limitation of its fixed functionality, certain domain-specific accelerators usually fail short on other types of operations as robotic algorithm evolving. Besides, the life span of such accelerators can be relatively short compared to the general-purpose approach since certain accelerators can be obsoleted fairly fast when they fail to keep up with the fast-evolving algorithms.
@@
From a high-level perspective, we observe that each architecture offers distinct advantages in terms of computational efficiency by leveraging various forms of parallelism: data-level, instruction-level, and thread-level—parallelism to achieve the compute efficiency required by specific types of robotic tasks with different area, power cost. Moving forward, our focus will be on three principal architectural categories: general-purpose CPUs, vector machines, and domain-specific accelerators. These architectures are strategically selected to meet the diverse computational needs of robotic systems within the stringent area and power constraints.
@@ III Methodology
This initial analysis in Section II allows us to narrow our focus to the following architecture categories: general-purpose CPUs, vector machines, and domain-specific accelerators. Those three types of architecture are fundamentally distinct in their designs, allowing a comparison of the following aspects:
  • General vs Specialized approach
  • Instruction set (CISC vs RISC)
  • Integration style (Tightly Integrated vs Decoupled)
  • Architectural spatial structure
Our investigation starts with contrasting general-purpose processors with specialized computational architectures. General-purpose processors includes CPU and vector extension on CPU support broad applicability across diverse tasks, whereas specialized domain-specific accelerators Gemmini are attributed to their highly efficient performance in specific computational scenarios.
The other distinct attribute of these architectures is their integration styles, ranging from standalone processing units (e.g. CPUs) to tightly integrated systems(e.g. CPU with vector machines) to loosely decoupled systems (e.g. CPU with Gemmini). This segment of the study highlights the significant performance enhancements and power efficiency gains achievable through strategic architectural integration. Saturn demonstrated a compact design of a vector machine tightly integrated into the CPU, allowing for precise interrupt and precise trap. On the other hand, Gemmini integrates with a RISC-V CPU using a decoupled RoCC interface, which provides the flexibility of allowing CPU and processing units to work concurrently. The reorder buffer in Gemmini decouples the instruction flow latency from the actual execution, allowing for latency hiding.
Another point of exploration in this comparison is to understand the impact of different instruction set architectures (ISAs). By comparing RISC-based systems like RISCV cores and Saturn against the CISC-style Gemmini, we aim to determine how the choice of ISA influences the overall architectural efficiency and application-specific performance. This comparison is pivotal in understanding the practical impacts of ISA designs on system behavior and performance metrics.
@@
To understand the tradeoff of the aspects mentioned above, we profiled representative architecture candidates from each categories17: For general-purpose CPUs, we evaluate RISC-V CPU cores, includes simple in-order core Rocket [1], superscalar in-order core Shuttle[20], superscalar out-of-order BOOM core[3], and more performant out-of-order cores such as MediumBOOM, LargeBOOM and MegaBOOM core with different numbers of FPUs[29], etc. For vector machines, we access Saturn, which is a vector extension on RISC-V CPU with short-vector microarchitecture [28]. For domain-specific accelerators, we profiled the CISC-style systolic array Gemmini[9].
@@ IV Workload Characterization
Robots require a diverse range of algorithms and workloads to function properly and efficiently. These workloads include reasoning, sensing, perception, localization, forward and inverse dynamics, mapping, motion planning and control, etc. Those algorithms are usually computationally intensive since they usually involves a wide variety of dense linear algebra kernels, such as general matrix-vector multiplication (GEMV) and general matrix-matrix multiplication (GEMM), and non-typical domain-specific operations, such as Cholesky decomposition, Riccati recursion, etc. The interdependence and interleaving of those operations propose a need for highly efficient compute and efficient data movement and management to handle these tasks.
As robots evolve towards more complex rigid dynamics with multi-joint systems, this introduces additional challenges for control algorithms. In real-life safety-critical applications, control algorithms must run at higher frequencies to meet real-time deadlines. As the final stage in a robotic end-to-end pipeline, with its output directly executed by the actuator, control algorithms face even more critical challenges in meeting stringent application deadlines. Deterministic computation time and low latency are crucial for deterministic control in safety-critical applications such as obstacle avoidance navigation, where every millisecond of processing time can be critical.
Moreover, the kernels in classical robotic control workloads are relatively small compared to other artificial intelligence workloads that are dominated by DNN-based neural network workloads, like large language models (LLMs) and diffusion networks. In classical robotic workloads, the dimensions of operands involved in those linear algebra kernels are typically of the order of magnitude of 10. Using model predictive control (MPC) for UAV as an example, its operand matrix and vector sizes correspond to the state space and action space of the UAV, which are matrices of size of 12x4. This poses challenges on latency hiding and requires relatively high instruction throughput to keep the instruction window full or leverage instruction-level parallelism to maintain backend utilization, assuming the backend can complete such small kernels relatively fast.
IV-A Target Workload
This paper focuses on one representative robotic workload: model predictive control (MPC). MPC is a powerful tool for controlling highly dynamic robotic systems subject to complex constraints. Besides, MPC’s scalability is dramatic; MPC’s computation time grows cubically with the robot’s state space and linearly with the time horizon, making MPC not feasible for complex robots. Besides, MPC algorithms run at a frequency of 100-500Hz, leading to high computational requirements. This also requires the compute backends to manage data and execute operations with low latency.
Specifically, we explore the state-of-the-art Model Predictive Control (MPC), TinyMPC [17] as a target algorithm. TinyMPC performs optimization using the alternating direction method of multipliers (ADMM) method. The algorithm alternates between solving three sets of variables - the primal, slack, and dual - until convergence. Solving the linear system in the primal update is the most expensive part of each ADMM iteration.
By exploiting the MPC problem structure, this step becomes the solution to an LQR problem, where K and D are the optimal feedback gain and feed-forward term for a given timestep and are computed using Riccati recursion. Given a long enough horizon, the Riccati equations converge to the solution of the infinite horizon LQR problem, which is a single gain matrix K infinity and cost-to-go P infinity. We save on memory by caching just these two matrices instead of the entire horizon of Ks and Ps. Only the linear terms of the Riccati equation need to be updated as part of each ADMM iteration. Because we already calculated P and K, we can also pre-compute the matrix inverse and other matrix coefficients. The result is an algorithm whose most expensive operations are matrix-vector products, allowing TinyMPC to achieve a low memory footprint while performing fast online computations.

### v2

@@ Abstract
Empowering resource-limited robots to execute computationally intensive tasks such as locomotion and manipulation is challenging. This project provides a comprehensive design space exploration to determine optimal hardware computation architectures suitable for model-based control algorithms. We profile and optimize representative architectural designs across general-purpose scalar, vector processors, and specialized accelerators. Specifically, we compare scalar CPUs, vector machines, and domain-specialized accelerators using kernel-level benchmarks and end-to-end representative robotic workloads. Our exploration provides a quantitative performance, area, and utilization comparison and analyzes the trade-offs between these representative distinct architectural designs. We demonstrate that architectural modifications, software, and system optimization can alleviate bottlenecks and enhance utilization. Finally, we propose a code generation flow to simplify the engineering work for mapping robotic workloads to specialized architectures.
@@ I Introduction
Robots are increasingly being deployed across diverse fields such as package delivery, high-precision surgical assistance, and autonomous industrial robots for warehouse assistance. Robots run a broad spectrum of algorithms and workloads to function effectively, including reasoning, sensing, perception, localization, forward and inverse dynamics, mapping, motion planning, and control. As robots evolve towards more complex rigid dynamics with multi-joint systems, this introduces additional challenges for control algorithms.
Model predictive control (MPC) is a widely used approach for controlling highly dynamic robotic systems subject to complex constraints. However, its computation demand grows cubically with the robot’s state space and linearly with the prediction horizon. This challenge becomes insurmountable on embedded SoCs as robots evolve into highly dynamic multi-joint rigid body systems. Additionally, control algorithms must run at a frequency of 100-400Hz to meet real-time deadlines in safety-critical applications. Ensuring low-latency computation is crucial in safety-critical applications such as obstacle avoidance and navigation, where every millisecond of processing time can be critical.
Traditional microcontrollers and embedded SoCs either lack the computational capacity to meet such latency requirements or consume excessive energy thus becoming impractical for battery-powered edge devices. As robots become more complex and miniaturized, onboard resources, such as compute power, energy, memory storage, and bandwidth, become increasingly constrained. Furthermore, the combination of computation-demanding algorithms, real-time requirements, and resource constraints makes it increasingly impractical to rely on a single hardware architecture to meet all demands on an embedded SoC.
Recognizing the heterogeneity of modern SoCs, this report explores hardware architectures suitable for modern robotic systems by profiling MPC algorithms on embedded platforms as an example. Specifically, we conduct a comprehensive design space exploration across scalar, vector, and matrix architectures as depicted in Figure 1. Our exploration provides a quantitative performance, area, and utilization comparison and analyzes the trade-offs between these architectural designs. Finally, we propose a code generation flow to simplify the engineering work for mapping robotic workloads to specialized architectures.
@@ II Overview of Architectures
CPUs are the most commonly used general-purpose computing platforms. Since CPUs are designed to handle a broad range of kernels, cheap CPUs, especially microcontrollers, were traditionally used for simple rule-based robotic kernels such as proportional-integral-derivative (PID) control solvers. However, microcontrollers fail to meet the computational demand for real-time robotic tasks with scaled-up models and DNN-based algorithms. This is mainly due to microcontrollers being typically slow with a compute frequency of 200 MHz. One solution to maximize a CPU’s available compute is by exploiting the out-of-order and superscalar architecture to drive instruction-level parallelism. However, the performance margin comes with the cost of area overhead and power consumption. An out-of-order and superscalar CPU can achieve up to 100 GFLOP/Sec. However, its power efficiency is less than 1 GOP/J, which is usually not compatible with power-constrained embedded devices [24].
@@
Vector extensions allow CPUs with lightweight front-ends to achieve similar performance as out-of-order and superscalar CPUs, while dramatically simplifying implementation by eliminating excess control logic. Vector machines exploit data parallelism by processing multiple data elements with a single instruction without adding too much power and area overhead. Various vector architectures range from traditional long-vector machines, as presented in Cray-style vector machines [19] to Packed SIMD (P-SIMD) machines that utilize VLIW instruction encodings, [12], to contemporary novel short-vector machines [26]. These architectures adeptly support a broad array of operations, such as scalar multiplication, general vector multiplication (GEMV), and matrix-matrix multiplication (GEMM). While vector cores are highly flexible and can handle a wide range of computations, they may not be as efficient for GEMM as dedicated systolic arrays or other matrix-specific accelerators, especially in handling large, batched operations. This is because vector machines scale inadequately with a large datapath compared to a systolic array-style processor. As datapath length (DLEN) scales up to feed in data sufficiently to the increased number of lanes, the bandwidth of the register file must scale accordingly.
@@
GPUs are designed to drive data-level parallelism (DLP) and thread-level parallelism (TLP) with up to over 100 thousand cores running simultaneously, enabling massive parallelism. Only a few robotic workloads[5, 15] can be benefited from such massive parallelism. However, a typical GPU can reach up to 10 TOPS/Sec of performance. Its high throughput makes it suitable for tasks that involve heavy matrix computations, graphics rendering, and other non-real-time robotic workloads. GPUs mainly exhibit TLP by running thousands of threads simultaneously to handle extensive graphical calculations and data processing tasks efficiently. However, a conventional GPU’s die area is typically around 600 mm2 to over 800 mm2 [7], which consumes too much area to be placed on micro- or nano-robotics where the system has limited space on a chip and faces weight constraints. In terms of energy consumption, GPUs consume 100W-350W of power[25], which is far beyond the power budget of most resource-limited robotic systems.
@@
In addition to the general-purpose processors, application-specific integrated circuits (ASICs), and field-programmable gate arrays (FPGAs) have recently been proposed as feasible platforms for power-constraint robotic applications since they offer higher energy efficiency than CPUs and GPUs. Both ASICs and FPGAs consume relatively less power and are often integrated into small systems with limited memory. Exemplified by their reconfigurability and programmability, FPGAs adapt well to the involving robotic algorithms compared to ASICs. Partial Reconfiguration (PR) extends this flexibility, allowing dynamic reconfiguration of a portion of the FPGA at runtime[22]. However, the highly specialized programming model, challenges in mapping irregular algorithms with control flow, and high resource overhead from configuration/routing logic make maximizing the utilization of all processing elements (PEs) challenging[23].
@@
With the integration of domain-specific accelerators, ASICs, and FPGAs can parallelize computations while leveraging the specific characteristics of the kernels to eliminate excess logic and simplify implementation[22] and offload specific computational tasks to the accelerators. Research [23, 13, 16] has been devoted to developing hardware-aware algorithms for FPGAs and has been proposed to achieve software-equivalent performance, making ASICs and FPGAs capable of meeting real-time requirements [14, 20, 21].
@@
Particularly, spatial array style domain-specific accelerators have been attributed to their energy efficiency and compute throughput in compute-intensive workloads such as neural networks. A typical spatial array accelerator architecture can achieve an efficiency of approximately 34.4 GOP/J while consuming 278 mW at a frame rate of 35 FPS for the AlexNet model[4]. The benefit of re-purposing deep neural network (DNN) accelerators for robot computing is based on the observation that neural network accelerators are optimized for matrix linear algebra [9, 11] while major robotic algorithms also mainly consist of DNN workloads and small-sized GEMM and GEMV [8, 2]. Most robotics chips already contain an ASIC neural net inference accelerator to speed up DNN-based workloads like perception, vision, and online learning. Suppose it can be shown that the same accelerator is also capable of speeding up robot dynamics computations. In that case, the overall SoC architecture can benefit from significant savings in the chip area, make room for other specialized units, or instead enable a physical design that is more cost-effective for ASIC implementation.
@@
Besides DNN accelerators, research efforts[20, 13, 5, 15, 16, 14, 21, 17] have been dedicated to design domain-specific robotic accelerators exclusively for robotic algorithms, focusing on a particular kernel defined for optimization problems for specific types of robots. Although domain-specific accelerators can offer improved efficiency over general-purpose processors on specific robotics workloads, the dramatic amount of engineering cost due to its specific programmable interface with its dedicated domain-specific language (DSL); Besides, due to the limitation of its fixed functionality, certain domain-specific accelerators usually fail short on other types of operations as robotic algorithm evolving. Besides, the life span of such accelerators can be relatively short compared to the general-purpose approach since certain accelerators can be obsoleted fairly fast when they fail to keep up with the fast-evolving algorithms.
@@
From a high-level perspective, we observe that each architecture offers distinct advantages in terms of computational efficiency by leveraging various forms of parallelism: data-level, instruction-level, and thread-level parallelism to achieve the compute efficiency required by specific types of robotic tasks with different area and power costs. Moving forward, our focus will be on three principal architectural categories: general-purpose CPUs, vector machines, and domain-specific accelerators. These architectures are strategically selected to meet the diverse computational needs of robotic systems within stringent area and power constraints.
@@ III Methodology
This initial analysis in Section II allows us to narrow our focus to the following architecture categories: general-purpose CPUs, vector machines, and domain-specific accelerators. To understand the tradeoff of these architectures, we profile representative architecture candidates from each categories1: For general-purpose CPUs, we evaluate RISC-V CPU cores, including simple in-order core Rocket [1], superscalar in-order core Shuttle[18], superscalar out-of-order BOOM core[3], and more performant out-of-order cores such as MediumBOOM, LargeBOOM and MegaBOOM core with different numbers of floating point unit (FPUs)[27]. For vector machines, we assess Saturn, a vector extension on RISC-V CPU with short-vector microarchitecture [26]. For domain-specific accelerators, we profile a state-of-the-art systolic array Gemmini[9].
One distinct attribute of these architectures is their integration styles, ranging from standalone processing units (e.g. CPUs) to tightly integrated systems (e.g. a CPU with vector extensions) to loosely coupled systems (e.g. a CPU with Gemmini). This segment of the study highlights the significant performance enhancements through software optimization and architectural changes. Saturn is a compact design of a vector machine tightly integrated into the CPU. On the other hand, Gemmini integrates with a RISC-V CPU using a decoupled Rocket custom coprocessor (RoCC) interface, which provides the flexibility of allowing CPU and processing units to work concurrently. The reorder buffer in Gemmini decouples the instruction flow latency from the actual execution, allowing for latency hiding.
@@ IV Workload Characterization
MPC algorithms are computationally intensive. MPC incorporates dense linear algebra kernels like GEMV and GEMM, and domain-specific operations such as Cholesky decomposition and Riccati recursion, which further increase the computational load. Specifically, we profile TinyMPC, a state-of-the-art embedded MPC algorithm as an example, and measure the impact of hardware architectures on kernel performance and end-to-end workloads.
IV-A Workload Structure
One state-of-the-art approach to solving MPC is the alternating direction method of multipliers (ADMM). ADMM alternates between updating the primal variables, slack variables, and dual variables, ensuring convergence to an optimal solution by gradually reducing constraint violations; with each subroutine denoted as primal update, dual update, slack update. TinyMPC redefines the most computationally expensive step, the primal update, as a linear quadratic regulator (LQR) problem. LQR equations are solved using Riccati recursion, where the feedback gain matrix $K$ and feedforward term $d$ are computed for each time step. Over sufficiently long horizons, the Riccati recursion converges to an infinite-horizon solution, represented by a single gain matrix $K_{\infty}$ and cost-to-go matrix $P_{\infty}$. TinyMPC optimizes memory use by caching these two matrices, avoiding the need to store the entire horizon of intermediate $K$ and $P$ matrices. During each ADMM iteration, only the linear terms of the Riccati equation need updating, further accelerating the computation. Pre-computed matrix inverses and coefficients ensure the algorithm’s most intensive operations are limited to matrix-vector products, enabling fast and memory-efficient execution suitable for constrained environments.